greciacisneros:

a weather balloon full of helium with a volume of 12.9 L at 22oc is sitting inside an aircraft hangar. the temperature of the hangar decreases to 13oc. which equation would be easiest to use to find the new volume of the balloon?

Vocaloid:

next time please post the answer choices A. P1V1 = P2V2 B. PV = nRT C. (V1/T1) = (V2/T2) D. P = F/A that being said, think about what you are given: initial volume, initial temperature, final temperature and what you need to solve for: final volume which equation will let you solve for the unknown?